### v2

# Self-eXplainable AI for Medical Image Analysis: A Survey and New Outlooks

Abstract
The increasing demand for transparent and reliable models, particularly in high-stakes decision-making areas such as medical image analysis, has led to the emergence of eXplainable Artificial Intelligence (XAI). Post-hoc XAI techniques, which aim to explain black-box models after training, have raised concerns about their fidelity to model predictions. In contrast, Self-eXplainable AI (S-XAI) offers a compelling alternative by incorporating explainability directly into the training process of deep learning models. This approach allows models to generate inherent explanations that are closely aligned with their internal decision-making processes, enhancing transparency and supporting the trustworthiness, robustness, and accountability of AI systems in real-world medical applications. To facilitate the development of S-XAI methods for medical image analysis, this survey presents a comprehensive review across various image modalities and clinical applications. It covers more than 200 papers from three key perspectives: 1) input explainability through the integration of explainable feature engineering and knowledge graph, 2) model explainability via attention-based learning, concept-based learning, and prototype-based learning, and 3) output explainability by providing textual and counterfactual explanations. This paper also outlines desired characteristics of explainability and evaluation methods for assessing explanation quality, while discussing major challenges and future research directions in developing S-XAI for medical image analysis.
{IEEEkeywords}
Self-eXplainable Artificial Intelligence (S-XAI), Medical Image Analysis, Input Explainability, Model Explainability, Output Explainability, S-XAI Evaluation
1 Introduction
Artificial intelligence (AI), particularly deep learning, has driven significant advancements in medical image analysis, including applications in disease diagnosis, lesion segmentation, medical report generation (MRG), and visual question answering (VQA). Deep neural networks (DNNs) automatically learn features from input data and produce optimal outputs. However, the inherent complexity nature of DNNs hinder our understanding of the decision-making processes behind these models. Consequently, DNNs are often considered as black-box models, which has raised concerns about their transparency, interpretability, and accountability for their successful deployment in real-world clinical applications [1].
Figure 1: Illustration of post-hoc XAI versus Self-eXplainable AI (S-XAI).
To tackle the challenge of developing more trustworthy AI systems, research efforts are increasingly focusing on various eXplainable AI (XAI) methods, enhancing transparency [2], fairness [3], and robustness [4]. However, most XAI methods aim to generate explanations for the outputs of black-box AI models after they have been trained, a category known as post-hoc XAI, as illustrated in Fig. 1 top. These methods utilize additional explanation models or algorithms to provide insights into the decision-making process of the primary AI model. In the field of medical image analysis, commonly used post-hoc XAI techniques include feature attribution methods, such as gradient-based approaches (e.g., LRP [5], CAM [6]) and perturbation-based approaches (e.g., LIME [7], Kernel SHAP [8]). Additionally, some methods explored concept attributions, learning human-defined concepts from the internal activations of DNNs (e.g., TCAV [9], CAR [10]). Post-hoc XAI techniques are often model-agnostic, indicating that they can be flexibly applied to a variety of already-trained black-box AI models.
Since post-hoc explanations are generated separately from the primary AI model, several valid concerns have been raised: 1) these explanations may not always be faithful to the actual decision-making process of black-box models [11, 12]; 2) they may lack sufficient detail to fully elucidate the model’s functioning [13]. These limitations of post-hoc XAI approaches are particularly problematic in high-stakes domains like medical image analysis, where clinicians require a deep and trustworthy understanding of how an AI model arrives at its predictions. The issues about the faithfulness and sufficiency of post-hoc explanations highlight the importance of exploring self-explainable AI models as a potentially more reliable and transparent alternative.
Self-eXplainable AI (S-XAI) is a category of XAI methods designed to be interpretable by nature, as illustrated in Fig. 1 bottom. These methods incorporate explainability as an integral part of the model during the training process, rather than generating explanations after the model has been trained. Conventional inherently interpretable methods include various white-box machine learning models, such as decision trees [14], generalized additive models [15], and rule-based systems [16]. In this survey, we focus primarily on DNNs and extend the characteristics of self-explainability across the entire pipeline, from model input to architecture to output, enabling direct inspection and understanding of the reasoning behind the model’s predictions without reliance on external explanation methods. In contrast to post-hoc XAI approaches, S-XAI methods aim to provide explanations that are inherent, transparent, and faithful, aligning directly with the model’s internal decision-making mechanisms. Such explanations are essential for the effective adoption and clinical integration of AI-powered decision support systems. Furthermore, S-XAI facilitates collaborative decision-making between clinicians and AI systems, fostering better-informed and more accountable medical diagnoses and interventions.
Figure 2: The upward trend of the total number of S-XAI research papers from 2018 to 2024 (Jan-Sept), where medical articles account for half.
This paper presents the first systematic review of S-XAI for medical image analysis, covering methodology, medical applications, and evaluation metrics, while also offering an in-depth discussion on challenges and future directions. Although there is a wealth of literature on medical XAI surveys [17, 2, 18, 19, 20, 21, 22] that deliver valuable insights, none have focused specifically on a comprehensive review of S-XAI methods applied to medical image analysis. We analyze more than 200 papers published from 2018 to October 2024, sourced from the proceedings of NeurIPS, ICLR, ICML, AAAI, CVPR, ICCV, ECCV, and MICCAI as well as top-tier journals in the field, including Nature Medicine, Transactions on Pattern Analysis and Machine Intelligence, Transactions on Medical Imaging, Medical Image Analysis, or those cited in related works. The statistics of research articles using keywords self-explainable, medical, image on Google Scholar are presented in Fig. 2, which reveal two key observations: 1) there has been a significant and consistent increase in research papers focused on self-explainable AI over the years, indicating growing interest and emphasis within the research community; 2) nearly half of the total research papers (green bars) are dedicated to applying S-XAI techniques in medical imaging (blue bars), highlighting the vital importance of S-XAI in the medical field.
To summarize, this survey presents insights into S-XAI for medical image analysis, with our contributions outlined below:
  1. 1.
    Novel scope of XAI survey: As an emerging XAI method that actively offers explainability from the model itself, S-XAI is attracting growing attention from the research community. This work serves as the first comprehensive survey on this topic.
  2. 2.
    Systematic review of methods: We introduce a novel taxonomy of relevant papers and review them based on input explainability, model explainability, and output explainability. This offers insights into potential technical innovations for S-XAI methods.
  3. 3.
    Thorough overview of applications: We overview various downstream applications across different anatomical locations and modalities in current S-XAI research. This illustrates the ongoing development of S-XAI technologies in medical image analysis, serving as a reference for future applications in diverse contexts.
  4. 4.
    Comprehensive survey of evaluations: We analyze a range of desired characteristics and evaluation methods to assess the quality of explainability. This provides guidelines for developing clinically explainable AI systems that are trustworthy and meaningful for end-users.
  5. 5.
    In-depth discussion of challenges and future work: We discuss the key challenges and look forward to the promising future directions. This underscores existing shortcomings and identifies new outlooks for researchers to drive further advancements.
2 S-XAI in medical image analysis
Figure 3: Overview of Self-eXplainable AI (S-XAI) frameworks including input explainability, model explainability, and output explainability.
Transparency and trustworthiness are essential for deep learning models deployed in real-world applications of medical image analysis. To address this need, the research community has explored various XAI methods and proposed several XAI taxonomy. According to existing literature [23, 17, 19], XAI methods can be categorized by the following criteria.
1) Intrinsic versus Post-hoc: This criteria differentiates whether interpretability is inherent to the model’s architecture (intrinsic) or achieved after the model training (post-hoc).
2) Model-specific versus Model-agnostic: Model-specific methods are restricted to particular model classes, whereas model-agnostic methods can be applied to explain any model;
3) Local versus Global: The scope of an explanation distinguishes between those for an individual prediction (local) or those for the entire model behavior (global).
4) Explanation modality: The common types of explanation include visual explanation, textual explanation, concept explanation, sample explanation, etc.
This survey concentrates on S-XAI methods for medical image analysis that allow models to inherently explain their own decision-making. As depicted in Fig. 3, we introduce a new taxonomy of S-XAI based on the three key components of DNNs.
1) Input explainability (Sec. 3): Input explainability focuses on integrating additional explainable inputs with deep features of medical images obtained from various anatomical locations and modalities to produce final predictions. By incorporating external knowledge and context-specific information, the accuracy and reliability of these predictions can be significantly improved.
2) Model explainability (Sec. 4): Model explainability aims to design inherently intepretable model architectures of DNNs. Instead of explaining a black-box model, transforming the model into an interpretable format enhances understanding of how it processes information.
3) Output explainability (Sec. 5): Output explainability refers to the model’s ability to generate not just predictions for various medical image tasks but also accompanying explanations through an explanation generator. This capability aids in understanding the rationale behind the model’s predictions, facilitating informed medical decision-making.
The following sections summarize and categorize the most relevant works on S-XAI methods applied to medical image analysis. Comprehensive lists of the reviewed S-XAI methods are provided, detailing the employed S-XAI techniques, publication year, anatomical location, image modality, medical application, and the datasets used.
3 Input Explainability
In this section, we will explore input explainability by integrating external domain knowledge, focusing on two key approaches, i.e., a) explainable feature engineering (Sec. 3.1) and b) knowledge graph (Sec. 3.2). As shown in Fig. 4, these explainable inputs will interact with the deep features of image inputs and be combined to support final predictions.
3.1 Explainable Feature Engineering
Feature engineering aims to transform raw images into a more useful set of human-interpretable features [24], which is essential for traditional machine learning. However, deep learning models automatically extract features from raw images, simplifying the handcraft process but often sacrificing interpretability. To enhance input explainability, it is beneficial to combine explainable feature engineering with deep learning, as shown in Fig. 4A. This integration enhances models’ interpretability by injecting domain knowledge into models and ensuring that the learned features are relevant and meaningful for clinical applications.
A common strategy is to combine both deep and handcrafted features from an input image to make final predictions [25, 26]. For example, Kapse et al.[25] introduce a self-interpretable multiple instance learning framework that simultaneously learns from deep image features and handcrafted morphometric and spatial descriptors. Local and global interpretability are demonstrated through quantitative and qualitative benchmarking. Another line of approach involves incorporating interpretable clinical variables as additional inputs alongside images, often utilizing multimodal learning techniques [27, 28]. For instance, Xiang et al.[27] introduce OvcaFinder, an interpretable model that combines deep learning predictions from ultrasound images with Ovarian–Adnexal Reporting and Data System scores provided by radiologists, as well as routine clinical variables for diagnosing ovarian cancer. This approach enhances diagnostic accuracy and explains the impact of key features and regions on the prediction outcomes.
Discussion: Although explainable feature engineering can be time-consuming, it brings valuable prior knowledge and enhances the interpretability of deep learning models concerning input features. Despite the increasing research in this area, most studies prioritize accuracy improvements, with limited analysis given to the explainability. Additionally, effective information fusion and interaction poses a key challenge.
3.2 Knowledge Graph
Figure 4: Input explainability that incorporates explainable feature engineering and knowledge graphs as additional inputs.
A knowledge graph (KG) is a structured representation of factual knowledge that captures relationships between entities in the real world [29]. Recently, there has been increasing interest in integrating structured domain knowledge into downstream applications [30, 31, 32], with the recognition that leveraging domain knowledge can greatly improve the performance and self-explainability of deep learning models. Regarding medical image analysis, the types of KG can be broadly categorized into three categories: 1) prior KG, which serves as a foundational resource that gathers existing domain expertise; 2) data KG, which is derived from the analysis of large-scale medical image datasets; and 3) hybrid KG, which combines prior KG and data KG, as illustrated in Fig. 4B.
3.2.1 Prior knowledge graph
In the medical domain, prior KG captures and organizes factual information about medical concepts and their relationships, which can be constructed from multiple sources, such as medical literature, medical ontologies, clinical guidelines, and expert opinions. It serves as a comprehensive repository of medical knowledge, encompassing details about diseases, symptoms, treatments, medications, anatomical structures, and provides a vital foundation for medical decision-making, clinical research, and healthcare analytics [33, 34, 35]. By harnessing the medical prior knowledge encoded in the graph, AI models introduce specialized domain information during the input phase and leverage this prior knowledge to enhance model’s self-explainability, gain valuable insights, and improve the performance of AI model in clinical task such as Med-VQA [36, 37] and medical report generation [38, 39, 40, 41]. For example, Zhang et al.[38] construct prior graphs from clinical studies to support radiology report generation. Huang et al.[40] develop symptom graphs based on a professional perspective of medical images, injecting the model with the ability to understand these images, thereby generating reliable medical reports. Liu et al.[41] extract a set of 52.6K triplets containing medical knowledge from OwnThink (https://www.ownthink.com) and use this external information to create SLAKE, a knowledge-enhanced bilingual dataset for training and testing Med-VQA systems. Prior KGs enhance S-XAI models by integrating medical facts. However, the creation of these KGs largely depends on specialized expertise, which makes the process labor-intensive and time-consuming. Furthermore, they often lack the adaptability required for analyzing dynamic clinical datasets.
3.2.2 Data knowledge graph
Unlike prior KG, which relies on expert insights, data KG is built directly from the dataset itself, giving them the potential to discover previously unknown relationships and correlations [42, 43, 44]. There are two main types of data KG for enhancing the explainability of AI models. 1) Intra-data KG is a local relation graph within a data sample [45, 46, 47, 48, 49]. For instance, Chen et al.[45] propose label co-occurrence learning to explore the relationship between pathologies in multi-label chest X-ray image classification. Huang et al.[48] construct the medical KG based on the types of diseases and questions from a Med-VQA dataset and propose a medical knowledge-based VQA network. 2) Inter-data KG captures relationships between different data samples [50, 51]. For example, Zheng et al.[50] utilize six meta-paths to connect four types of data (i.e., X-ray, CT, ultrasound, and text descriptions of diagnosis) and create a multi-modal KG for COVID-19 diagnosis. Furthermore, Zhao et al.[52] and Qi et al.[53] leverage both intra-data and inter-data graph into a unified framework for disease diagnosis and localization in X-rays. Constructing a data KG involves leveraging the inherent characteristics of the dataset itself to assist S-XAI models. However, these methods would harbor inherent biases that can vary significantly across different datasets.
3.2.3 Hybrid knowledge graph
Hybrid KG integrates prior KG and data KG, which not only provides a foundation of established medical facts, but also allows for the integration of new information and the refinement of existing knowledge. Consequently, hybrid KG combines strengths of both prior KG and data KG, offering a more comprehensive and adaptable knowledge representation for S-XAI models in the medical field [54, 55, 56, 57, 58, 59, 60]. For instance, Zhou et al.[54] combine intra- and inter-contrastive attentions to learn abnormal attended visual features of X-rays and then leverage a chest radiology prior KG to enhance the thoracic disease diagnosis models. To enable dynamic graph construction, Li et al.[57] initially utilize a foundational structure from prior KG and then add nodes or modify their relationships based on the specific knowledge relevant to each X-ray image. Moreover, Hu et al.[60] use large language models to extract labels and build a large-scale medical VQA dataset, and then leverage graph neural networks to learn logical reasoning paths.
Discussion: The utilization of medical KGs for S-XAI poses both challenges and promising opportunities. First, integrating diverse prior medical knowledge into a graph format is labor-intensive and costly, requiring constant updates and refinements to incorporate the latest knowledge. Another challenge lies in the heterogeneity of medical data. The expanding variety of data modalities would complicate the effective integration within KGs. Developing robust AI algorithms to extract meaningful features from medical data and align them with medical KGs remains an ongoing research endeavor.
4 Model Explainability
In this section, we present model explainability by designing interpretable model architectures, such as attention-based learning (Sec. 4.1), concept-based learning (Sec. 4.2), and prototype-based learning (Sec. 4.3).
4.1 Attention-based Learning
Attention-based learning aims to capture specific image regions that are relevant to specific model task while suppressing irrelevant regions based on attention maps [61]. Therefore, attention-based models can provide visual explanations that interpret model decision-making [19, 2]. We categorize attention-based S-XAI models into 1) structure-guided attention models and 2) loss-guided attention models. As illustrated in Fig. 5, the former estimates attention scores through structural design, whereas the latter constrains attention maps using loss functions.
4.1.1 Structure-guided attention models
Figure 5: Attention-based learning, including 1) structure-guided and 2) loss-guided attention models. X-ray images borrowed from [62].
As shown in the top branch of Fig. 5, structure-guided attention models are characterized by the attention estimator whose outputs directly influence model’s predictions. The estimator structure combine the features from different encoder-layer features to calculate feature attention scores as the generated attention map to effectively explain the model’s predictions [63, 64, 65]. For example, Jetley et al. [63] first introduce attention-based learning for XAI by proposing an attention estimator to calculate feature compatibility scores and weigh feature maps, which are then used directly as input for a linear classifier. Numerous studies have incorporated multiple attention mechanisms for medical image classification and segmentation tasks [66, 67, 68, 69, 65, 70]. For example, Schempler et al. [67] extend the attention estimator by extracting local information from coarse features for attention gates, facilitating more fine-grained visual interpretation for lesion segmentation and ultrasound diagnosis. Similarly, Gu et al. [69] develop a comprehensive attention module with spatial, channel, and scale attention. The segmentation results on skin lesions and fetal organs demonstrate improved performance and better interpretability of target area positioning and scale. Beyond 2D data, how to use attention to explain more complex 3D medical image analysis is more challenging. Lozupone et al. [70] present an attention module that fuses attention weights from sagittal, coronal, and axial slices to diagnose Alzheimer’s disease on 3D MRI brain scans, in which 3D attention maps can be visualized to explain the model’s decision-making process. For fast MRI reconstruction, Huang et al. [71] propose a shifted windows deformable attention mechanism which uses reference points to impose spatial constraints on attention and combines the attention outputs from different windows to produce reconstruction results. Although structure-guided attention models can provide explanations for predictions, they are still difficult to align with clear human-understandable decision-making basis.
4.1.2 Loss-guided attention models
As shown in the bottom branch of Fig. 5, loss-guided attention models use interpretable labels (i.e., reference attention maps) to constrain generated attention maps from attention mechanism by loss functions. Using ground-truth masks of regions of interest (RoIs) to guide the generation of attention maps is a widely adopted approach in medical image classification [72, 73, 74, 75, 62]. For instance, Yang et al. [72] directly optimize the attention maps by a Dice loss, which encourages the model to focus on target areas that are highly relevant to the classification of breast cancer microscopy images. To alleviate the challenge of obtaining pixel-level annotations, Yin et al. [75] pre-train a histological feature extractor to identify significant clinically relevant feature masks, which are then used to guide and regularize attention maps. By considering the varying contributions of histological features for classification, the model can selectively focus on different features based on the distribution of nuclei in each instance. In medical image segmentation, labels corresponding to edges and shapes of specific regions are often reused to guide attention in learning semantic information [76, 77, 78]. Sun et al.[76] combine spatial attention with the attention estimator in U-Net decoders, enabling the model to interpret learned features at each resolution. They also introduce a gated shape stream alongside the texture stream, where the shape attention maps are aligned with actual edges by a binary cross-entropy loss, enhancing cardiac MRI segmentation.
Compared with lesion masks, eye tracking data provides a more accurate depiction of expert focus, as it captures the way doctors visually process information during diagnosis. Bhattacharya et al. [62] leverage the captured doctors’ attention to guide model training. They employ a teacher-student network to replicate the visual cognitive behavior of doctors when diagnosing diseases on chest radiographs. The teacher model is trained based on the visual search patterns of radiologists, and the student model utilizes an attention loss to predict attention from the teacher network using eye tracking data.
Discussion: Attention-based S-XAI methods offer effective visual explanations. Structure-guided attention models use attention-weighted outputs as inputs for predictors, reflecting decision-making but often lacking clear semantic information. In contrast, loss-guided attention models provide explicit semantic details in their explanations, though their attention outputs do not directly influence model decisions. Overall, while these methods improve transparency and insights into decision-making, the understandability of attention maps and their relevance still require further investigation.
4.2 Concept-based Learning
Concept-based S-XAI methods provide explanations in terms of high-level, human-interpretable attributes rather than low-level, non-interpretable features, enabling users to gain deeper insights into the underlying reasoning with easily understandable concepts [9, 79]. It also helps in identifying model biases and allows for adjustments to enhance performance and trustworthiness. Most concept-based S-XAI methods focus on making decisions based on a set of concepts while also detailing the contribution of each concept to the final prediction [80, 81, 82, 83]. We propose to categorize concept-based S-XAI methods into three types: 1) supervised concept learning, 2) unsupervised automatic concept discovering, and 3) generative concept learning, as shown in Fig. 6.
The term Concept has been defined in different ways, which commonly represents high-level attributes [84, 9, 79]. In this paper, we suggest adopting a simple and clear categorization: Textual Concepts and Visual Concepts. Textual Concepts refer to textual descriptions of attributes associated with the classes. For example, in Fig. 6, the textual concepts for the classes (i.e., “pneumonia” and “normal”) include terms like Opacity, Effusion, Infiltration, etc. Visual Concepts, on the other hand, consist of semantically meaningful features within the image that may not be explicitly described in natural language. For example, Alvarez Melis et al. [4] consider the intermediate image representations in a space of interpretable atoms with low dimension as the concepts of a given image.
Figure 6: Concept-based learning, including 1) supervised concept learning, 2) unsupervised automatic concept discovering, and 3) generative concept learning.
4.2.1 Supervised concept learning
Supervised concept learning methods train deep learning models using annotations of textual concepts, particularly by supervising an intermediate layer to represent these concepts. A notable example is Concept Bottleneck Model (CBM) [80], which first maps latent image features to a concept bottleneck layer, where the number of neurons corresponds to the number of human-defined concepts, and then predicts final results based on the concept scores from this layer. Learning concept representations supervised by concept labels enables CBMs to directly show each concept’s contribution to the final prediction using the neuron values of the last layer. Additionally, CBMs allow model editing. When domain experts find certain predicted concept importance unreasonable, they can adjust the model’s predictions by modifying the weights of the concept bottleneck layer (i.e., test-time intervention). The CBM architecture has inspired many researchers to develop self-explainable methods, resulting in a series of CBM-like models [85, 86, 87].
The self-explainable nature of concept-based learning models has led to its application in medical imaging to satisfy the trustworthiness requirements of healthcare [88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98]. Chauhan et al. [88] propose Interactive CBMs, which can request labels for certain concepts from a human collaborator. This method exhibits good interpretability on chest and knee X-ray datasets. Kim et al. [95] present a medical concept retriever, which connects medical images with text and densely scores images on concept presence. This work uses a CBM architecture to develop an inherently interpretable model and conducts data auditing and model interpretation. Different from CBMs, Concept Whitening [99, 100] aims to whiten the latent space of neural networks and aligns the axes of the latent space with known concepts of interest. Zhao et al. [101] introduce a hybrid neuro-probabilistic reasoning algorithm for verifiable concept-based medical image diagnosis, which combines clinical concepts with a Bayesian network. However, a significant challenge in supervised concept learning is the scarcity of concept annotations, which require labor-intensive efforts from human experts. Therefore, some researchers prefer unsupervised automatic concept discovering, as it eliminates the need for extra annotations.
4.2.2 Unsupervised automatic concept discovering
These models modify their internal representations to identify visual concepts within image features, without relying on explicit annotations. The discovered concepts may not be directly associated with human-specified textual concepts, but can also be visualized and detailed their contributions to final predictions. Self-Explaining Neural Networks (SENN) [4] first utilize a concept encoder to extract clusters of image representations corresponding to different visual concepts, and adopt a relevance parametrizer to calculate the relevance scores of concepts. The final prediction is determined by the combination of discovered concepts and the corresponding relevance scores. Inspired by SENN, Sarkar et al. [102] utilize both a concept encoder and a concept decoder, which map images into concept space and use the concepts to reconstruct the original images, respectively. Since medical concept annotations are costly and require experts’ efforts, unsupervised automatic concept discovering is usually adopted to offer concept-based explanations in medical image analysis without expert-annotated labels. For example, Fang et al. [103] address the practical issue of classifying infections by proposing a visual concept mining method to explain fine-grained infectious keratitis images. They first use a saliency map based potential concept generator to discover visual concepts, and then propose a visual concept-enhanced framework that combines both image-level representations and the discovered concept features for classification. Although unsupervised automatic concept discovering can offer concept-based explanations, these explanations are abstract and usually cannot be directly described in natural language. To alleviate this issue while also addressing the lack of concept annotations, generative concept learning has become a promising research direction.
4.2.3 Generative concept learning
Leveraging foundation models, such as Large Language Models (LLMs) and Vision-Language Models (VLMs), can assist in generating and labeling textual concepts. A notable generative concept learning method, namely Language Guided Bottlenecks (LaBo) [104], employs an LLM (GPT-3 [105]) to generate textual concepts for each image category, which are filtered to form the concept bottleneck layer. LaBo then uses a pre-trained VLM (CLIP [106]) to calculate the similarity between input images and the generated concepts to obtain concept scores. The final prediction is based on the multiplication of a weight matrix and these concept scores. Label-free CBM [107] employs a similar pipeline, but trains an independent network that includes a concept bottleneck layer. In the medical domain, Kim et al. [108] enhance LaBo by incorporating a more fine-grained concept filtering mechanism and conduct explainability analysis on dermoscopic images, achieving performance improvements compared to the baseline. Similarly, some research [109, 110] employ ChatGPT and CLIP for explainable disease diagnosis. Bie et al. [111] propose an explainable prompt learning framework that leverages medical knowledge by aligning the semantics of images and clinical concept-driven prompts at multiple granularities, where category-wise clinical concepts are obtained by eliciting knowledge from LLMs.
Discussion: Concept-based learning hold significant importance in medical research and applications, particularly in advancing evidence-based medicine. The lack of fine-grained concept annotations and the performance-explainability trade-off are the limitations of concept-based methods. As there are more and more medical foundation models being developed, incorporating knowledge from these models and medical experts to efficiently annotate concept labels for datasets will be a promising and meaningful direction. Besides the most popular classification task, other medical applications of concept-based approaches should be further explored.
4.3 Prototype-based Learning
Figure 7: Prototype-based learning, including 1) explicit prototype and 2) implicit prototype. X-ray images borrowed from [112].
Prototype-based S-XAI models aims to provide a decision-making process where a model reasons through comparisons with a set of interpretable example prototypes. This reasoning aligns with human recognition patterns, as humans often identify objects by comparing them to example components [113]. These models first extract features from a given image and then compare the feature maps with the prototypes to calculate similarities. This process is considered interpretable because the decision making can be clearly attributed to the contribution of each interpretable prototype (e.g., by the similarity scores). According to how the prototypes are obtained, we define and categorize them to two types: 1) explicit prototype and 2) implicit prototype, as presented in Fig. 7. Explicit prototypes are specific high-dimensional feature representations extracted from certain training images, whereas implicit prototypes are latent high-dimensional representations that are close to a set of typical images’ representations. All existing prototype-based S-XAI models do not require supervision at the prototype level and aim to automatically find meaningful prototypes to facilitate interpretable decision making.
4.3.1 Explicit prototype based models
The first model of this type is ProtopNet [114], which introduces a three-stage training scheme: 1) Freeze the final layer, and only train the feature extractor. 2) Replace the learned representations in the prototype layer with the nearest feature patch from the training set. 3) Remain the feature extractor fixed and fine-tune the parameters of the final layer. Later works closely follow this training scheme while addressing different limitations of this initial framework [115, 116, 117, 118, 119, 120, 121, 122]. Adopting prototype-based S-XAI models in the medical domain presents additional challenges. Unlike natural images where the representative prototype occupies an area with a relatively stable size, medical image features such as disease regions in chest X-ray images can vary significantly in size. To address this, XProtoNet [112] proposes to predict an occurrence map and summing the similarity scores within those areas, rather than relying solely on the maximum similarity score as done in ProtopNet. Similarly, [123] introduces prototypes with square and rectangular spatial dimensions for COVID-19 detection in chest X-rays. In evaluations of ProtopNet, Mohammadjafari et al.[124] observe a performance drop for Alzheimer’s disease detection using MRI, whereas Carlon et al.[125] report a high-level of interpretability satisfaction from radiologists in breast mass classification using mammograms. In mammogram based breast cancer diagnosis, Wang et al.[126] propose to leverage knowledge distillation to improve model performance. To overcome the confounding issue in mammogram based mass lesion classification, Barnett et al.[127] employ a multi-stage framework that identifies the mass margin features for malignancy prediction, skipping image patches that have already been used in previous prototypes during the prototype projection step to improve prototype diversity. In brain tumor classification, MProtoNet [128] introduces a new attention module with soft masking and online-CAM loss applied in 3D multi-parametric MRI. To predict the brain age based on MR and ultrasound images, Hesse et al.[129] utilize the weighted mean of prototype labels. Additionally, INSightR-Net [130] formulates the diabetic retinopathy grading as a regression task and apply the prototype based framework, while ProtoAL [131] explores an active learning setting for prototype-based models in diabetic retinopathy.
Although these models offer interpretability in a one-to-one mapping to the input image, they can also make it difficult for users to identify which specific property is important in the corresponding image patch (e.g., is it the color or texture that matters in this prototypical area?). This issue can be partially mitigated using implicit prototype based models.
4.3.2 Implicit prototype based models
This type of model follows a similar training scheme as the models based on explicit prototypes, with the major difference in avoiding the prototype replacement step, or only projecting the prototype to the training images’ feature patches for visualizations. This scheme is simpler than one that includes prototype replacement step and has different interpretability benefits. Li et al.[132] propose the earliest work using latent prototypes, which leverages a decoder to visualize the meanings of the learned prototypes. Protoeval [133] designs a set of loss functions to encourage the learned latent prototypes to be more stable and consistent across different images. Recently, to address the concern that prototype-based models often underperform their black box counterparts, Tan et al.[134] develop an automatic prototype discovery and refinement strategy to decompose the parameters of the trained classification head and thus guarantees the performance. Although a serious of implicit prototype based models [135, 136, 137] are demonstrated effectiveness on natural images, their developments and applications on medical image analysis need to be further explored.
Figure 8: Output explainability that provides (a) textual explanations, including fully-structured, semi-structured, and free-structured text; and (b) counterfactual explanations. The difference (red box) indicates the explanation. X-ray images borrowed from [138].
Discussion: In terms of performance, implicit prototype based models generally outperform explicit ones, probably due to the greater flexibility in prototype learning. Regarding interpretability, both types of models offer unique advantages. Explicit prototypes can be intuitively explained through one-to-one mappings to the input image, while implicit prototypes can be explained using a diverse set of images with similar activations. However, current prototype-based S-XAI models in medical image analysis primarily use explicit prototypes, making the exploration of implicit prototypes in this domain a promising area for future research.
5 Output Explainability
This section discusses output explainability by generating explanations alongside model predictions, including textual (Sec. 5.1) and counterfactual (Sec. 5.2) explanations.
5.1 Textual Explanation
S-XAI models that provide textual explanations generate human-readable descriptions to accompany their predictions as part of outputs, similar to image captioning. These methods use natural language to clarify model decisions and typically require textual descriptions for supervision. Some studies explore the integration of textual explanations with visual ones. We categorize these methods into three types based on the structure of textual explanations: 1) fully-structured, 2) semi-structured, and 3) free-structured text, as shown in Fig. 8(a).
5.1.1 Fully-structured text generation
To address the challenges posed by complex unstructured medical reports, early efforts transformed target texts into fully structured formats, such as descriptive tags [139], attributes [140, 141] with fixed templates, rather than natural language. For example, Medical Subject Headings (MeSH) [142] can be utilized to describe image content instead of unstructured reports. Therefore, both Shin et al.[143] and Gasimova et al.[144] employ CNN-RNN frameworks to identify diseases and generate corresponding MeSH sequences, detailing location, severity, and affected organs in chest X-ray images. However, complete descriptions in natural language are more human-understandable than a set of simple tags, leading several studies to focus on generating reports in a semi-structured format.
5.1.2 Semi-structured text generation
Generating semi-structured text involves a partially structured format with predefined topics and constraints. For instance, pathology report generation methods [145, 146, 147] produce reports that focus on describing certain types of cell attributes along with a concluding statement. Additionally, Wang et al.[148] introduce a hierarchical framework for medical image explanation, which first predicts semantically related topics and then incorporates these topics as constraints for the language generation model. More recently, some studies have focused on generating individual sentences based on anatomical regions [149, 150, 151, 152]. For example, Tanida et al.[151] introduce a Region-Guided Radiology Report Generation method that identifies unique anatomical regions in the chest and generates specific descriptions for the most salient areas, ensuring each sentence in the report is linked to a particular anatomical region. Overall, semi-structured approaches effectively balance the rigidity of fully structured reports with the inconsistency of completely free-text reports.
5.1.3 Free-structured text generation
With the advancement of language models, reports generated for a given input image are no longer limited to structured formats; instead, they now focus on more open, free-structured text descriptions. These approaches typically combine an image encoder to extract visual features with a language model to produce coherent sentences [153]. Several research efforts provide comprehensive explanations that include both textual and visual justifications for diagnostic decisions [154, 155, 156, 157].
In addition to directly generating textual explanations, some research studies have incorporated the classification of pathological terms or tags in two distinct ways. The first approach utilizes a “classification-report generation” pipeline, integrating a classifier within the report generation network to enhance feature representations [158, 159]. For example, Yuan et al.[158] further employ a sentence-level attention mechanism alongside a word-level attention model to analyze multi-view chest X-rays, using predicted medical concepts to improve the accuracy of medical reports. Conversely, the second approach follows a “report generation-classification” pipeline, leveraging interpretable region-of-interest characterization for final diagnoses. For instance, Zhang et al.[160] construct a pathologist-level interpretable diagnostic framework that first detects tumour regions in whole slide images (WSIs), then generates natural language descriptions of microscopic findings with feature-aware visual attention, and finally establishes a diagnostic conclusion. Moreover, integrating region localization and lesion segmentation can enhance the quality of textual explanations [161, 162, 163, 164]. For instance, Wang et al.[161] develop a Text-Image Embedding network that incorporates multi-level attention to highlight meaningful text words and X-ray image regions for disease detection and reporting.
The emergence of LLMs and VLMs offers a more interactive and comprehensible method for generating textual explanations. Recent medical VLMs applied to various medical images, such as chest X-rays (e.g., XrayGPT [165]), skin images (e.g., SkinGPT [166]), and general medical images (e.g., Med-flamingo [167], LLaMa-Med [168], MedDr [169], HuatuoGPT-Vision [170]), can analyze and respond to open-ended questions about the input images, thanks to their pretraining on extensive datasets of image-report pairs. Take XrayGPT [165] as an example, given an input image, this combined model can address open-ended questions, such as “What are the main findings and impressions from the given X-ray?”. These models not only excel in report generation, but also demonstrate exceptional capability in delivering comprehensive explanations for a wide range of medical inquiries.
Discussion: Textual explanations have demonstrated significant effectiveness in providing human-interpretable judgments through natural language. This type of S-XAI approach has become especially valuable with the advancement of language models, enabling the generation of lengthy reports and the ability to answer open-ended questions. However, it is crucial to enhance the quality and reliability of these generated textual explanations. Some recent studies utilize techniques such as knowledge decoupling [171] and instruction tuning [172] to address challenges like hallucination, thereby improving the effectiveness and trustworthiness of textual explanations.
5.2 Counterfactual Explanation
Counterfactual explanations describe a causal situation by imagining a hypothetical reality that contradicts the observed facts: If X had not occurred, Y would not have occurred [23]. These explanations present a contrastive example: for a given image, its counterfactual image can alter the model’s prediction to a predefined output through the minimal perturbation to observations on the original image, as illustrated in the diagram of Fig. 8(b). Traditional counterfactual explanations are generated through a post-hoc paradigm [173], that is, a classification model is first trained as a black-box model, and then a generative model such as GAN [174] is applied to produce the counterfactual counterpart. However, post-hoc counterfactual explanations are susceptible to issues related to the classifier’s robustness and complexity (e.g., overfitting and excessive generalization), resulting in explanations that are inadequate for effective interpretability [175].
To tackle this issue, some research has investigated self-explainable variants of counterfactual explanations. An alternative is to incorporate a generative model directly into the predictor, training them jointly so that the model can generate explanations for its own predictions. For example, CounterNet [176] combines the training of the predictive model with the generation of counterfactual explanations in an end-to-end framework. Compared to post-hoc approaches, it is able to produce counterfactuals with higher validity. Similarly, VCNet et al.[177] utilizes a counterfactual generator based on a conditional variational autoencoder, allowing for control and adjustments in the latent space to create more realistic counterfactuals. In the medical field, Wilms et al.[178] introduce an invertible, self-explainable generative model based on efficient normalizing flow for brain age regression and brain sex classification on 3D neuroimaging data. This model can generate predictions during the forward process and produce explanations, including voxel-level attribution maps and counterfactual images, in the reverse process.
Discussion: S-XAI models that generate counterfactual explanations alongside predictions show greater promise than post-hoc methods. However, their application in medical image analysis remains largely underexplored.
6 Evaluation
Table 1: Desirable characteristics of explanations and explainability methods.
Type Characteristic Description Ref.
Faithfulness, Fidelity, Truthfulness Explanations should truthfully reflect the AI model decision process. [179, 180, 181]
Consistency, Invariance, Robustness For a fixed model, explanation of similar data points (with similar prediction outputs) should be similar. [181, 7, 182]
      Explanations Understandability, Comprehensibility Explanations should be easily understandable by clinical users without requiring technical knowledge. [181, 179, 182]
Clinical Relevance Explanation should be relevant to physicians’ clinical decision-making pattern, and can support their clinical reasoning process. [181]
Plausibility, Factuality, Persuasiveness Users’ judgment on explanation plausibility (i.e., how convincing the explanations are to humans) may inform users about AI decision quality, including potential flaws or biases. [179, 181, 7]
Explainability Methods Computational Complexity The computational complexity of explanation algorithms. [183, 181]
Generalizability, Portability To increase the utility because of the diversity of model architectures. [183]
Assessing explainability presents significant challenges. In this section, we will outline the desired characteristics and evaluation methods for explainability.
6.1 Desired Characteristics of Explainability
It is important for S-XAI models to possess certain desirable qualities when providing explanations. In medical applications, the characteristics of high-quality explanations should align closely with the real-world necessities of clinical practice. Van et al.[19] and Adadi et al.[184] summarize several essential traits of XAI methods for medical image analysis and healthcare, respectively. Jin et al.[181] propose five criteria for optimizing clinical XAI. With the development of LLMs, usability has emerged as a key factor that enhances a model’s credibility [185], with interactive and dynamic explanations being preferred over static ones. Table 1 presents expected traits based on our literature review regarding explanations and explainability methods.
6.2 Evaluation Methods
Doshi-Velez and Kim [186] propose three distinct categories for evaluating XAI methods. 1) Application-grounded evaluations engage experts specific to a field to evaluation explanation quality. 2) Human-grounded evaluations involve non-experts assessing the overall quality of explanations. 3) Functionality-grounded evaluations use proxy tasks instead of human. In the medical field, it is crucial to involve domain experts in the evaluation process, ideally in contexts that utilize real tasks and data [2].
6.2.1 Human-centered evaluation
Conducting human-centered evaluations with medical experts is crucial to determine user satisfaction with the explanations provided by S-XAI models. These evaluations can assess explanation quality using both qualitative and quantitative metrics.
Qualitative metrics include evaluating the usefulness, satisfaction, confidence, and trust in provided explanations through interviews or questionnaires [187]. For instance, the System Causability Scale [188] can be used to measure the quality of interpretability methods applicable in the medical field. Gale et al.[141] assess experts’ acceptance of explanations by scoring each type on a 10-level Likert scale.
Quantitative metrics focus on measuring task performance of human-machine collaboration with factors such as accuracy, response time, likelihood of deviation, ability to detect errors, and even physiological responses [187]. For example, Sayres et al.[189] demonstrate that AI assistance improves diagnostic accuracy, subjective confidence, and time spent, based on investigations involving ten ophthalmologists with varying experience levels in grading diabetic retinopathy.
Overall, human-centered evaluations offer the significant advantage of providing direct and compelling evidence of the effectiveness of explanations [186]. However, they can be costly and time-consuming, as they require recruiting expert participants and obtaining necessary approvals. Most importantly, these evaluations are inherently subjective.
6.2.2 Functionality-grounded evaluation
This category of evaluation can be employed to assess the fidelity of explanations. The accuracy of S-XAI methods in generating genuine explanations is referred to as the fidelity of an explainer. We will present a variety of functionality-grounded evaluation methods for different types of explanations.
Attention-based explanations can be assessed through causal metrics in the absence of references, such as deletion and insertion [190] and RemOve And Retrain (ROAR) [191]. They can also be evaluated by comparing with ground truth annotations such as semantic masks [73, 192] or human expert eye fixation [193]. Additionally, the Activation Precision metric [194] is introduced to quantify the proportion of relevant information from the relevant region used to classify the mass margin based on radiologist annotations.
Concept-based explanations can primarily be evaluated using metrics Concept Error [80, 102], T-CAV score [9], Completeness Score [195], and Concept Relevance [4, 84]. Other evaluation methods include Concept Alignment Score, Mutual Information [85], and Concept Purity [119].
Prototype-based explanations can be measured by non-representativeness (i.e., how well the examples represent the explanations) and diversity (i.e., the degree of integration within the explanation) [196]. Additionally, a consistency score and a stability score are also used to assess prototype-based explanations [133].
Textual explanations assessments involve using metrics such as BLEU [197], ROUGE-L [198], and CIDEr [199] to compare generated natural language descriptions against ground truth reference sentences provided by experts. A benchmark study [17] of textual explanations is conducted on chest X-ray image datasets.
Counterfactual explanations can be evaluated by counterfactual validity, proximity, sparsity, and diversity [200]. Other metrics include Frechet Inception Distance score, Foreign Object Preservation score, as well as clinical metrics to illustrate the clinical utility of explanations [201].
7 Challenges and Future Directions
Despite the rapid advancements in S-XAI for medical image analysis, several significant challenges remain unresolved. In this section, we will analyze the existing challenges and discuss potential future directions to enhance the effectiveness and reliability of S-XAI in the medical domain.
7.1 Benchmark Construction
Establishing benchmarks for S-XAI in medical image analysis is essential, which enables standardize evaluations and fair comparisons between different methods.
7.1.1 Dataset
A major challenge of dataset construction is the limited availability of doctors to annotate large-scale medical images, especially in S-XAI, which requires fine-grained annotations such as concepts and textual descriptions. Current medical datasets that meet interpretability standards often have limited volumes, reducing the generalizability and applicability of S-XAI methods in real-world contexts.
7.1.2 Evaluation
Automated evaluation of explanations poses another significant challenge. In the medical field, human-centered evaluations often depend on clinician expertise, but variability in expert opinions can lead to biased and subjective assessments [202]. Additionally, existing functionality-grounded evaluations still rely on manual annotations.
To tackle these challenges, future directions include leveraging semi-automated annotation tools to assist clinicians in the annotation process, thereby reducing their workload. Furthermore, developing objective metrics and standardized protocols to assess the quality of model explanations will be a critical research trend in S-XAI.
7.2 S-XAI in the Era of Foundation Models
Medical foundation models, such as LLMs [203, 204, 205] and VLMs [206, 165, 167], are trained on extensive medical corpora, allowing them to inherently encode rich domain-specific knowledge. The convergence of S-XAI with these large models creates substantial opportunities for developing trustworthy medical AI systems in the future [185].
7.2.1 S-XAI benefits foundation models
Foundation models are typically large, with an enormous number of parameters, making it challenging to explore their decision-making processes. This complexity can lead to potential biases and a lack of transparency. In addition to using post-hoc techniques [207, 208] to interpret foundation models, S-XAI methods can enhance input explainability through explainable prompts [209] and knowledge-enhanced prompts [210].
7.2.2 Foundation models advance S-XAI
Foundation models learn useful representations from the clinical knowledge in medical corpora [211]. By leveraging their advanced capabilities, S-XAI methods can generate user-friendly explanations [212] and enable flexible generative concept-based learning [213]. Additionally, foundation models can aid in evaluating S-XAI methods that mimic human cognitive processes [214].
7.3 S-XAI with Human-in-the-loop
Integrating Human-in-the-loop (HITL) processes is crucial for effectively implementing S-XAI in the medical field. This approach not only enhances the overall performance of AI systems but also fosters trust among medical experts.
7.3.1 Improving accuracy through human intervention
A HITL framework allows for the identification and removal of potential confounding factors, such as artifacts or biases in datasets. For instance, clinicians can adjust predicted concepts, leading to a more accurate concept bottleneck model [90]. This collaborative approach significantly enhances the prediction accuracy by incorporating expert insights.
7.3.2 Enhancing explainability through human feedback
To ensure continuous improvement, a versioning or feedback evaluation system should be established, to build trust during hospital evaluations. This requires fostering collaboration between S-XAI researchers and clinical practitioners to systematically gather and utilize feedback for model refinement.
A challenge in integrating HITL processes is the variability in clinician expertise and availability, which can impact the consistency and quality of feedback. Ensuring effective integration of human knowledge into the AI training process without introducing biases or errors is a complex task.
7.4 Trade-off between Performance and Interpretability
It is commonly believed that as model complexity increases to boost performance, interpretability tends to decline [215, 216]. Conversely, more interpretable models may sacrifice prediction accuracy. However, some researchers argue that there is no scientific evidence for a general trade-off between accuracy and interpretability [217]. Recent advancements in concept-based models [100, 95, 96] have shown performance comparable to black-box models in medical image analysis. This success relies on the ability to identify patterns in an interpretable way while accurately fitting the data [13]. Future S-XAI methods are expected to optimize both performance and interpretability, potentially establishing a theoretical foundation for this balance.
7.5 Other Explainability of S-XAI
7.5.1 Multi-modal explainability
Medical data often exists in various forms, such as images, texts, and omics. By integrating these modalities, multi-modal S-XAI methods can provide more comprehensive explanations that align with clinicians’ processes. Additionally, S-XAI methods can be used to identify correlations during multi-modal data fusion [218], offering significant potential for discovering new biomarkers.
7.5.2 Causal explainability
Another direction for S-XAI focuses on causality, which defines the cause-and-effect relationship and can be mathematically modeled [219]. Traditional deep learning methods in medical imaging often confuse correlation with causation, leading to potentially harmful errors. For instance, DeGrave et al.[220] found that COVID-19 diagnosis methods are primarily identifying spurious correlations using XAI techniques. Several efforts [221, 222] have been made to address dataset biases, which would be highly beneficial.
8 Conclusion
This survey reviews recent advancements in self-explainable artificial intelligence (S-XAI) for medical image analysis. Contrary to previous surveys that primarily focus on post-hoc XAI techniques, this paper emphasizes inherently interpretable S-XAI models, which are gaining traction in research. This survey introduce S-XAI from three key perspectives, i.e., input explainability, model explainability, and output explainability. Additionally, this survey explore the desired characteristics of explainability and various evaluation methods for assessing explanation quality. While significant progress has been made, it also highlights key challenges that need to be tackled and provides insights for future research on trustworthy AI systems in clinical practice. Overall, this survey serves as a valuable reference for the XAI community, particularly within the medical imaging field, and lays the groundwork for future advancements that will improve the transparency and trustworthiness of AI tools in healthcare.
.1 Lists of reviewed S-XAI methods
We provide comprehensive lists of our reviewed S-XAI methods for medical image analysis, including knowledge graph (Table 2), attention-based learning (Table 3), concept-based learning (Table 4), prototype-based learning (Table 5), and textual explanation (Table 6). Each table contains the information of the S-XAI technique employed, publication year, anatomical location, image modality, medical application, and the used datasets.
Table 2: Input explainability methods based on knowledge graph (KG). The abbreviations here are CLS: classification, DET: detection, MRG: medical report generation, VQA: visual question answering.
Method Year Location Modality Task Dataset KG Type
Naseem et al.[36] 2023 Multiple Pathology VQA  [223] Prior KG
Guo et al.[37] 2022 Multiple X-ray, CT, MRI VQA  [41] Prior KG
Zhang et al.[38] 2020 Chest X-ray MRG  [224] Prior KG
Liu et al.[39] 2021 Chest X-ray MRG  [225][226] Prior KG
Huang et al.[40] 2023 Chest X-ray MRG  [225][226] Prior KG
Liu et al.[41] 2021 Multiple X-ray, CT, MRI VQA  [41] Prior KG
Chen et al.[45] 2020 Chest X-ray CLS [224, 227] Data KG
Hou et al.[46] 2021 Chest X-ray CLS  [225][226] Data KG
Liu et al.[47] 2021 Chest X-ray MRG  [225][226] Data KG
Huang et al.[48] 2023 Multiple X-ray, CT, MRI, US VQA [41] Data KG
Li et al.[49] 2024 Multiple Pathology CLS [228] Data KG
Zheng et al.[50] 2021 Chest X-ray, CT, US, text CLS private Data KG
Liu et al.[51] 2021 Breast Mammogram DET  [229] Data KG
Zhao et al.[52] 2021 Chest X-ray DET  [224] Data KG
Qi et al.[53] 2022 Chest X-ray DET  [224] Data KG
Zhou et al.[54] 2021 Chest X-ray CLS  [224, 227] Hybrid KG
Wu et al. [55] 2023 Chest X-ray CLS  [226] Hybrid KG
Li et al.[56] 2019 Chest X-ray MRG  [225] Hybrid KG
Li et al.[57] 2023 Chest X-ray MRG  [225][226] Hybrid KG
Kale et al.[58] 2023 Chest X-ray MRG  [225] Hybrid KG
Hu et al.[59] 2023 Chest X-ray VQA  [59] Hybrid KG
Hu et al.[60] 2024 Chest X-ray VQA  [59] Hybrid KG
Table 3: Model explainability methods based on attention-based learning. The abbreviations here are CLS: classification, SEG: segmentation, IRE: image reconstruction, REG: regression.
Method Year Location Modality Task Dataset Attention Type
Li et al. [65] 2021 Eye Retinal images CLS [230] Structure-Guided
Wang et al. [66] 2018 Breast X-ray CLS [231] Structure-Guided
Schempler et al. [67] 2019 Abdominal, Fetal CT, US SEG + DET [232, 233] Structure-Guided
Lian et al. [68] 2019 Brain MRI REG [234],[235] Structure-Guided
Gu et al. [69] 2020 Skin, Fetal Dermatology, MRI SEG [236] Structure-Guided
Lozupone et al. [70] 2024 Brain MRI CLS [234] Structure-Guided
Huang et al. [71] 2022 Head MRI IRE [237] Structure-Guided
Bhattacharya et al. [62] 2022 Chest X-ray CLS [238, 239, 240, 241, 242, 243, 244, 245, 246, 247] Loss-Guided
Yang et al. [72] 2019 Breast Histopathology CLS [248] Loss-Guided
Yan et al. [73] 2019 Skin Dermatology CLS [249], [250] Loss-Guided
Barata et al. [74] 2021 Skin Dermatology CLS [250], [236] Loss-Guided
Yin et al. [75] 2021 Liver Histopathology CLS [251] Loss-Guided
Sun et al. [76] 2020 Cardiac MRI SEG [252, 253] Loss-Guided
Karri et al. [77] 2022 Skin, Brain, Abdominal Dermatology, MRI, CT SEG [254, 255, 256] Loss-Guided
Li et al. [78] 2023 Skin Dermatology SEG [250, 236, 257, 254] Loss-Guided
Table 4: Model explainability methods based on concept-based learning. The abbreviations here are CLS: classification.
Method Year Location Modality Task Dataset Concept
Koh et al. [80] 2020 Knee X-ray CLS [258] Supervised (CBM)
Chauhan et al. [88] 2023 Knee, Chest X-ray CLS [258][227] Supervised (CBM)
Patricio et al. [89] 2023 Skin Dermatology CLS [259, 260] Supervised (CBM)
Yan et al. [90] 2023 Skin Dermatology CLS Private Supervised (CBM)
Bie et al. [91] 2024 Skin Dermatology CLS [259, 260, 261] Supervised (CBM)
Marcinkevivcs et al. [92] 2024 Appendix US CLS Private Supervised (CBM)
Lucieri et al. [93] 2022 Skin Dermatology CLS [259], [260], [257] Supervised
Jalaboi et al. [94] 2023 Skin Dermatology CLS  [262], [263] Supervised
Kim et al. [95] 2024 Skin Dermatology CLS [257, 264, 265, 261, 259] Supervised (CBM)
Pang et al. [96] 2024 Skin, Blood cell Dermatology, Microscopy CLS [261], [266] Supervised (CBM)
Wen et al. [97] 2024 Eye Retinal images CLS [267], [268] Supervised
Gao et al. [98] 2024 Multiple Dermatology, Pathology, US, X-ray CLS [254, 269, 270, 271, 226] Supervised
Hou et al. [100] 2024 Skin Dermatology CLS [259], [261] Supervised
Zhao et al. [101] 2021 Chest CT CLS [272] Supervised
Fang et al. [103] 2020 Eye Slit lamp microscopy CLS [273] Concept discovery
Kim et al. [108] 2023 Skin Dermatology CLS [254] Generated concept
Liu et al. [109] 2023 Multiple X-ray, CT CLS [239, 274, 270] Generated concept
Patricio et al. [110] 2024 Skin Dermatology CLS [259], [260], [236] Generated concept
Bie et al. [111] 2024 Multiple Dermatology, X-ray CLS [259, 261, 275, 239] Generated concept
Table 5: Model explainability methods based on prototype-based learning. The abbreviations here are CLS: classification, REG: regression.
Method Year Location Modality Task Dataset Prototype Type
Kim et al.[112] 2020 Chest X-ray CLS [224] Explicit
Singh et al.[123] 2021 Chest X-ray CLS [276] Explicit
Mohammadjafari et al.[124] 2021 Brain MRI CLS [277] Explicit
Carloni et al.[125] 2022 Breast Mammogram CLS [278] Explicit
Wang et al.[126] 2022 Breast Mammogram CLS [279] Explicit
Barnett et al.[127] 2021 Breast Mammogram CLS Private Explicit
Wei et al.[128] 2024 Brain MRI CLS [280] Explicit
Hesse et al.[129] 2024 Brain MRI, US REG [281], [282] Explicit
Hesse et al.[130] 2022 Eye Retinal images REG [283] Explicit
Santos et al.[131] 2024 Eye Retinal images CLS [284] Explicit
Table 6: Output explainability methods that provide textual explanations. The abbreviations here are MRG: medical report generation, CLS: classification, LOC: location, SEG: segmentation, VQA: visual question answering, Vis: Visual explanation.
Method Year Location Modality Task Dataset Text Type Vis.
Pino et al.[139] 2021 Chest X-ray MRG [225], [226] Fully-structured $\checkmark$
Rodin et al.[140] 2019 Chest X-ray MRG [226] Fully-structured $\checkmark$
Gale et al.[141] 2019 Pelvic X-ray MRG Private Fully-structured $\checkmark$
Shin et al.[143] 2016 Chest X-ray CLS + MRG [275] Fully-structured -
Gasimova et al.[144] 2019 Chest X-ray CLS + MRG [275] Fully-structured -
Zhang et al.[145] 2017 Bladder Pathology MRG Private Semi-structured $\checkmark$
Zhang et al.[146] 2017 Bladder Pathology MRG Private Semi-structured $\checkmark$
Ma et al.[147] 2018 Cervix Pathology MRG Private Semi-structured $\checkmark$
Wang et al.[148] 2019 Chest X-ray CLS + MRG [225] Semi-structured $\checkmark$
Tanida et al.[151] 2023 Chest X-ray LOC + MRG [149] Semi-structured $\checkmark$
Wang et al.[152] 2022 Chest X-ray CLS + MRG [225], [226] Semi-structured -
Singh et al.[153] 2019 Chest X-ray MRG [225] Free-structured -
Spinks et al.[154] 2019 Chest X-ray MRG [225], [285] Free-structured $\checkmark$
Liu et al.[155] 2019 Chest X-ray MRG [225], [226] Free-structured $\checkmark$
Chen et al.[156] 2020 Chest X-ray MRG [225], [226] Free-structured $\checkmark$
Wang et al.[157] 2023 Chest X-ray MRG [225], [226] Free-structured $\checkmark$
Yuan et al.[158] 2019 Chest X-ray CLS + MRG [227, 225] Free-structured $\checkmark$
Lee et al.[159] 2019 Breast Mammogram CLS + MRG [229] Free-structured $\checkmark$
Zhang et al.[160] 2019 Bladder Pathology CLS + MRG [228] Free-structured $\checkmark$
Wang et al.[161] 2018 Chest X-ray LOC + MRG [224], [275] Free-structured $\checkmark$
Jing et al.[162] 2018 Multiple X-ray, Pathology LOC + MRG [225], [286] Free-structured $\checkmark$
Zeng et al.[163] 2020 Multiple US, X-ray LOC + MRG [275] Free-structured $\checkmark$
Tian et al.[164] 2018 Abdomen CT SEG + MRG  [287] Free-structured $\checkmark$
Thawkar et al.[165] 2023 Chest X-ray VQA [226] Free-structured -
Zhou et al.[166] 2024 Skin Dermatology VQA [261, 288] Free-structured -
Moor et al.[167] 2023 Multiple Multiple VQA [289] Free-structured -
Li et al.[168] 2024 Multiple Multiple VQA  [290, 41, 223] Free-structured -
He et al.[169] 2024 Multiple Multiple VQA  [290, 223, 226] Free-structured -
Chent et al.[170] 2024 Multiple Multiple VQA  [290, 41, 223, 291] Free-structured -
Kang et al.[171] 2024 Chest X-ray VQA [225, 292, 226] Free-structured -
Chen et al.[172] 2024 Chest X-ray VQA  [172] Free-structured -
.2 Public datasets used in S-XAI
We provide an overview of more than 70 datasets currently available for S-XAI in the medical image domain. Table LABEL:tab:all_dataset presents the key characteristics of these datasets, including modality, scale, and task. In this section, we will introduce the relevant datasets categorized by image modalities, highlighting their contributions to the development of S-XAI. For more detailed information about these datasets, we direct readers to the relevant publications and sources.
.2.1 Radiology
Radiological images generally include modalities such as X-ray, MRI, CT, mammography, and ultrasound. Among these, X-ray is one of the most commonly used modalities in S-XAI for medical image analysis. For instance, the SLAKE [41] dataset collects knowledge triplets from the open source knowledge graph to assist the model in achieving a better understanding of X-ray images. Medical-CXR-VQA [60] focuses on the five types of questions (i.e., abnormality, presence, view, location and type) and aligns them with the key information of the X-ray image, resulting more reliable answers. VQA-RAD [290] is manually constructed based on clinicians asking naturally occurring questions about radiology images and providing reference answers, resulting in a dataset rich in quality expertise and knowledge to capture the details of radiology images. OAI [258] provides knee X-rays for knee osteoarthritis grading and offers clinical concepts (e.g., joint space narrowing, bone spurs, calcification), making it suitable for concept-based learning. Moreover, datasets such as IU X-ray [225], MIMIC-CXR [226], OpenI [275], and Chest ImaGenome [293] provide a large number of chest X-rays along with corresponding free-text reports, which can facilitate the generation of textual explanations. Finally, MIMIC-CXR-VQA [292], CheXbench [172], SLAKE [41], and MIMIC-Diff-VQA [59] are constructed for the VQA task, providing support for interactive explanations of S-XAI models.
Regarding MRI, SUN09 [252] and AC17 [253] provide cardiac segmentation masks, while BraTS 2020 [255] offers brain masks. However, most MRI datasets have a limited number of samples, which may affect the generalization ability of models. For CT images, datasets like CT-150 [232], NIH-TCIA CT-82 [233], and LiTS [287] can be used for segmentation tasks, while LIDC-IDRI [272] provides lung cancer annotations with grade of eight attributes, benefiting concept-based learning. Additionally, disease diagnoses on mammography datasets [278, 229, 294] and ultrasound datasets [271, 282] also serve as applications for S-XAI methods.
.2.2 Dermatology
In the scope of dermatology, datasets with fine-grained concept annotations are commonly used in concept-based learning for S-XAI. Derm7pt [259] is a dermoscopic image dataset containing 1,011 images with clinical concepts for melanoma skin lesions according to the 7-point checklist criteria [295]. PH2 dataset [260] includes 200 dermoscopic images of melanocytic lesions with segmentation masks and several clinical concepts. SkinCon [261] is a skin disease dataset containing 3,230 images with 48 clinical concepts densely annotated by dermatologists for fine-grained model debugging and analysis, where the images are selected from the Fitzpatrick 17k [264] and DDI [265] skin image datasets. Other datasets such as ISIC [250, 236, 257] and HAM10000 [254] are also broadly used datasets but without explicit fine-grained concept annotations. Dermoscopic image datasets significantly facilitate the development of S-XAI, however, annotating concept labels requires the efforts of human experts and is labor-intensive. Hence there are currently only a few datasets with a limited number of samples that have fine-grained concept labels. This poses a significant challenge for concept-based S-XAI, especially in supervised concept learning.
.2.3 Pathology
With regard to pathological image datasets, PathVQA [223] is the first dataset focused on pathology VQA, featuring over 32K open-ended questions derived from 4,998 pathology images. PEIR Gross [286] contains 7,442 image-caption pairs across 21 different sub-categories, with each caption consisting of a single sentence. The Cancer Genome Atlas (TCGA) [228] provides multimodal data for over 20,000 tumor and normal samples, offers multimodal data for more than 20K tumor and normal samples, encompassing clinical data, DNA, and various imaging types (diagnostic images, tissue images, and radiological images). Additionally, datasets such as BACH [248], Biopsy4Grading [251], and NCT [269] are available for classifying breast cancer, non-alcoholic fatty liver disease, and colorectal cancer, respectively. WBCAtt [266] provides 113K microscopic images of white blood cells, annotated with 11 morphological attributes categorized into four main groups: overall cell, nucleus, cytoplasm, and granules. Since pathological images are the “gold standard” for cancer diagnosis, the development of S-XAI models in pathology is highly significant.
.2.4 Retinal images
Regarding retinal disease classification, EyePACS [283] is a large-scale dataset for grading diabetic retinopathy, containing more than 88K images. ACRIMA [230] offers 705 images for glaucoma assessment. In addition to classification labels, FGADR [267], DDR [268], and IDRID [270] also provide fine-grained masks for segmenting various types of lesions.
.2.5 Others
Other medical datasets utilized in S-XAI include the Hyperkvasir endoscopy dataset [296] and the Infectious Keratitis slit lamp microscopy dataset [273]. Additionally, recent datasets like PMC-OA [289] and PMC-VQA [291] collect data from open-source medical literature corpus. These datasets encompass a wealth of multimodal data, which significantly facilitate the development of medical foundation models.
Table 7: Public datasets used in the reviewed S-XAI methods.
Dataset Modality Scale Task
SLAKE [41] X-ray 642 images, 14,028 QA pairs VQA, SEG, DET
ChestXray [224] X-ray 108,948 images CLS, MRG, DET, LOC
IU X-ray [225] X-ray 8,121 images, 3,996 texts MRG
MIMIC-Diff-VQA [59] X-ray 700K QA pairs, 164K images VQA
OAI [258] X-ray 26,626,000 images CLS
CheXbench [172] X-ray 6.1M QA pairs VQA
CheXpert [227] X-ray 224K images CLS
MIMIC-CXR [226] X-ray 377K images, 227K texts CLS, MRG
BCDR-F03 [231] X-ray 736 images CLS
RSNA [238] X-ray 30,000 images CLS, DET
ZhangLabData [239] OCT, X-ray 108,312 OCT images, 5,232 X-ray images CLS
SIIM-FISABIO-RSNA [240] X-ray 10,178 images CLS, DET
Public Radiography [241] X-ray 3,487 images CLS
COVQU [242] X-ray 18,479 images CLS, SEG
COVID-19-NY-SBU [243] X-ray, CT, MRI 1,384 cases CLS
MIDRC-RICORD-1C [244] X-ray 361 cases CLS
NIH [246] X-ray 108,948 images CLS
VinBigData [247] X-ray 18,000 images CLS, DET
Montgomery [274] X-ray 138 images CLS
COVID-19 [276] X-ray 761 images CLS
OpenI [275] X-ray 7,470 images, 3,955 texts CLS, MRG
Chest ImaGenome [149] X-ray 242K images, 217K texts CLS, MRG, DET
MIMIC-CXR-VQA [292] X-ray 377K images with QA pairs VQA
ADNI-1 [234] MRI 818 subjects CLS, REG
ADNI-2 [235] MRI 599 subjects CLS, REG
SUN09 [252] MRI 395 slices SEG
AC17 [253] MRI 200 volumes SEG
BraTS 2020 [255] MRI 494 cases SEG, CLS
Calgary Campinas MRI [237] MRI 359 subjects SEG
OASIS [277] MRI 416 cases CLS
BraTS 2014 [280] MRI 65 scans SEG
IXI [281] MRI 600 cases REG
RICODR [245] CT, X-ray 240 CT scans, 1,000 X-ray images CLS
CT-150 [232] CT 150 scans SEG
Pancreas-CT [233] CT 82 scans SEG
CHAOS [256] CT, MRI 40 CT scans, 120 MRI scans SEG
LIDC-IDRI [272] CT 1,018 scans CLS, SEG
LiTS [287] CT 201 scans SEG
VQA-RAD [290] CT, MRI, X-ray 3,515 QA pairs, 315 images VQA
DDSM [229] Mammogram 2,620 cases CLS, MRG, SEG
CBIS-DDSM [278] Mammogram 1,644 cases CLS, SEG
CMMD [279] Mammogram 1,775 cases CLS
BUSI [271] Ultrasound 780 images CLS
FGLS [282] Ultrasound 4,290 volumes REG
HAM10000 [254] Dermatology 10,015 images CLS
ISIC 2016 [249] Dermatology 1,279  images CLS, SEG
ISIC 2017 [250] Dermatology 2,750 images CLS, SEG
ISIC 2018 [236] Dermatology 15,121 images CLS, SEG
ISIC 2019 [257] Dermatology 33,569 images CLS
Derm7pt [259] Dermatology 1,011 images CLS
PH2 [260] Dermatology 200 images CLS, SEG
SkinCon [261] Dermatology 3,230 images CLS
Fitzpatrick 17k [264] Dermatology 16,577 images CLS
DDI [265] Dermatology 656 images CLS
DermNetNZ [262] Dermatology 25K images CLS
SD-260 [263] Dermatology 6,584 images CLS
Dermnet [288] Dermatology 18,856 images CLS, VQA
PathVQA [223] Pathology 32K QA pairs, 4,998 images VQA
BACH [248] Histopathology 500 images CLS
Biopsy4Grading [251] Histopathology 351 images CLS
WBCAtt [266] Microscopy 113,278 images CLS
NCT [269] Histopathology 100K images CLS
TCGA [228] Pathology 20K studies CLS, MRG
PEIR Gross [286] Pathology 7,442 image-text pairs MRG
ACRIMA [230] Retinal images 705 images CLS
FGADR [267] Retinal images 2,842 images CLS, SEG
DDR [268] Retinal images 13,673 images CLS, SEG, DET
IDRID [270] Retinal images 516 images CLS, SEG, LOC
EyePACS [283] Retinal images 88,702 images CLS
Messidor [284] Retinal images 1,200 images CLS
Hyperkvasir [296] Endoscopy 110,079 images, 374 videos SEG, DET
Infectious Keratitis [273] Slit lamp microscopy 115,408 images CLS
PMC-OA [289] Multiple 1.65M image-text pairs VQA
PMC-VQA [291] Multiple 227K QA pairs, 149K images VQA